OpenStudy (anonymous):

What must be added to 3x -7 to make x^2 + 4x -1?

OpenStudy (zzr0ck3r):

to make \(3x\) got to \(4x\) we add \(x\) to make \(-7\) got to \(-1\) we add \(6\) to make \(0\) go to \(x^2\) we add \(x^2\) so in all we add \(x^2+x+6\) and we see \((3x-7)+(x^2+x+6)=3x-7+x^2+x+6=x^2+3x+x+6-7 = x^2+4x-1\)

OpenStudy (zzr0ck3r):

\((3x-7)+(x^2+x+6)=3x-7+x^2+x+6=x^2+3x+x+6-7\\ = x^2+4x-1\)

OpenStudy (zzr0ck3r):

sorry that got cut off...